# Rare and exclusive few-body decays of the Higgs, Z, W bosons, and the top quark

### v3

@@ III.2.3 Higgs decays into a gauge boson plus a flavoured meson
On the one hand, calculations for the inclusive flavour-changing $\rm H\to qq^{\prime}$ rates exist [110, 111, 91], but without the photon (or Z boson) emission and the exclusive final-state meson formation. These predicted FCNC Higgs branching fractions, $\mathcal{B}(\rm H\to qq^{\prime})\equiv\mathcal{B}(\rm H\to\mathrm{q}\overline{\mathrm{q}}^{\prime}+q^{\prime}\overline{q})$, amount to
@@
$\mathcal{B}(\rm H\to sb)$ $=$ $(8.9\pm 1.5)\cdot 10^{-8},\;\,\quad\mathcal{B}(\rm H\to ds)=1.19\cdot 10^{-11}.$ (17)
On the other hand, the work of [13] determined the $\rm H\to\gamma+VM$ branching fractions, and the work [103] provided the $\rm H\to Z+VM$ decay widths assuming (arbitrary) $\mathcal{O}(1)$ flavour-changing Yukawa couplings. From the results of [13] (resp., [103]), the branching ratios of Higgs decaying into a photon (resp., a Z boson) plus a flavoured neutral meson can be determined through the following EFT + LCDA-based expressions
@@
where $p,\ \lambda$ are the momentum and polarization of the vector meson, respectively, and $\mu$ is the factorization scale. The corresponding numerical values used here are [113, 114]: $f^{\perp}_{\rm B^{*0}}(3\text{ GeV})=0.2\text{ GeV}$, $f^{\perp}_{\rm B^{*0}_{\rm s}}(3\text{ GeV})=0.236\text{ GeV}$, $f^{\perp}_{\rm K^{*0}}(2\text{ GeV})=0.717f_{\rm K^{*0}}$, $f^{\perp}_{\rm D^{*0}}(2\text{ GeV})=0.202\text{ GeV}$. These transverse decay constants can be evolved to the higher scale $\mu_{\rm HZ}$ using the renormalization group equation [14], thereby giving: $f^{\perp}_{\rm VM}(\mu_{\rm HZ})=0.86\ f^{\perp}_{\rm VM}(3\text{ GeV})=0.83\ f^{\perp}_{\rm VM}(2\text{ GeV})$, using $\alpha_{\mathrm{s}}(\mu_{\rm HZ})=0.126$, $\alpha_{\mathrm{s}}(3\text{ GeV})=0.25$, $\alpha_{\mathrm{s}}(2\text{ GeV})=0.29$, and $N_{\mathrm{f}}=5$ number of flavours.
@@
[…]
@@
$\begin{cases}|\kappa_{\rm bs}|^{2}+|\kappa_{\rm sb}|^{2}=2.4\times 10^{-8},\\ |\kappa_{\rm bd}|^{2}+|\kappa_{\rm db}|^{2}=1.0\times 10^{-9},\\ |\kappa_{\rm cu}|^{2}+|\kappa_{\rm uc}|^{2}=7.4\times 10^{-21},\\ |\kappa_{\rm ds}|^{2}+|\kappa_{\rm sd}|^{2}=3.3\times 10^{-12}.\end{cases}$ (23)
Combining Eq. (23) with (18) or (19), we arrive at the theoretical branching fractions listed in Table 6. Such exclusive FCNC radiative decays are extremely suppressed in the SM, in the range of $\mathcal{O}(10^{-14}$$10^{-30})$. Therefore, radiative flavoured meson decays of the Higgs boson appear utterly rare to be visible at any current or future collider, and therefore a very clean probe of any BSM physics that may enhance Higgs FCNC decays. Experimental limits have been recently set for the $\rm H\to\gamma+K^{*0},\gamma+D^{*0}$ channels at $\mathcal{O}(10^{-5})$ [95], to be compared with our $\mathcal{O}(10^{-19})$ SM prediction.
Table 6: Compilation of exclusive Higgs decays to a gauge boson plus a flavoured meson. For each decay, we provide the branching fraction predicted using Eqs. (18), (19), as well as the current experimental upper limit and that estimated for HL-LHC. The last two columns indicate whether the decay can be produced at FCC-ee/FCC-hh.
@@
$\rm H\to\,X$ $+$ $\rm M$ Branching fraction Framework 2024 HL-LHC FCC-ee FCC-hh
$\rm H\,\to\,\gamma$ + $\rm K^{*0}$  $1.6$$\times~{}10^{-19}$ EFT+LCDA (this work) $<$ $2.2$×$\times~{}10^{-4}$ [95] $\lesssim 3.3\times~{}10^{-5}$
$\rm D^{*0}$  $6.7$$\times~{}10^{-27}$ EFT+LCDA (this work) $<$ $1.0$×$\times~{}10^{-3}$ [115] $\lesssim 1.5\times~{}10^{-4}$
@@
$\rm H\,\to\,Z$ + $\rm K^{*0}$  $1.4$$\times~{}10^{-21}$ EFT+LCDA (this work)
@@ III.2.4 Higgs decays into a W boson plus a meson
The charged $\rm H\to W^{\pm}M^{\mp}$ decays (corresponding to the diagrams of Fig. 4 with W bosons in the final state) differ qualitatively from the neutral radiative decays discussed above, because the W attaches itself to a charged current, and one can probe flavour-violating couplings of the Higgs boson. The theoretical complication is that the W has both transverse and longitudinal polarizations, yielding lengthier analytical expressions [12, 103]. Table 7 lists the theoretical predictions for Higgs decays into a W± boson plus a charged meson, and Fig. 8 presents them in graphical form. The rates for these processes range roughly between $10^{-5}$ and $10^{-10}$ with small differences between models for the same final state. The EFT + NRQM theoretical predictions of Ref. [10], which update the $\mathcal{B}$ numerical values computed in [12] (but have likely a typo in the $\mathcal{B}(\rm H\to B^{*\pm}W^{\mp})=10^{-5}$ rate quoted, which should be $10^{-10}$), agree with the alternative EFT + LCDA rates of Ref. [103]. There has been no experimental search to date at the LHC for any of these 11 decays. The most promising channel, with a $\mathcal{O}(10^{-5})$ branching fraction, is $\rm H\to W^{\pm}+\rho^{\mp}$ with the rho meson decaying into two pions with almost 100% probability, but seemingly experimental evidence for 7 (all) of them will only be possible at FCC-ee (FCC-hh).
Table 7: Compilation of exclusive Higgs decays to a W boson plus a meson. For each decay, we provide the predicted branching fraction(s) and the theoretical approach used to compute it. The last two columns indicate whether the decay can be produced at FCC-ee/FCC-hh.
@@
$\rm H\to\,W$ $+$ $\rm M$ Branching fraction Framework 2024 HL-LHC FCC-ee FCC-hh
$\rm H\,\to\,W^{\mp}$ + $\rm\pi^{\pm}$  $(4.2\pm 0.2)$$\times~{}10^{-6}$ EFT+NRQM [10]
 $(4.3\pm 0.2)$$\times~{}10^{-6}$ EFT+LCDA [103]
$\rm\rho^{\pm}$  $(1.5\pm 0.1)$$\times~{}10^{-5}$ EFT+NRQM [10]
 $(1.09\pm 0.05)$$\times~{}10^{-5}$ EFT+LCDA [103]
$\rm K^{\pm}$  $(3.3\pm 0.1)$$\times~{}10^{-7}$ EFT+NRQM [10]
 $(3.3\pm 0.1)$$\times~{}10^{-7}$ EFT+LCDA [103]
$\rm K^{*\pm}$  $(4.3\pm 0.2)$$\times~{}10^{-7}$ EFT+NRQM [10]
 $(5.6\pm 0.4)$$\times~{}10^{-7}$ EFT+LCDA [103]
$\rm D^{\pm}$  $(5.8\pm 0.6)$$\times~{}10^{-7}$ EFT+NRQM [10]
 $(5.6\pm 0.5)$$\times~{}10^{-7}$ EFT+LCDA [103]
$\rm D^{*\pm}$  $(1.3\pm 0.1)$$\times~{}10^{-6}$ EFT+NRQM [10]
 $(1.04\pm 0.14)$$\times~{}10^{-6}$ EFT+LCDA [103]
$\rm D_{s}^{\pm}$  $(1.6\pm 0.1)$$\times~{}10^{-5}$ EFT+NRQM [10]
 $(1.71\pm 0.11)$$\times~{}10^{-5}$ EFT+LCDA [103]
$\rm D_{s}^{*\pm}$  $(3.5\pm 0.2)$$\times~{}10^{-5}$ EFT+NRQM [10]
 $(2.51\pm 0.19)$$\times~{}10^{-5}$ EFT+LCDA [103]
$\rm B^{\pm}$  $(1.6\pm 0.4)$$\times~{}10^{-10}$ EFT+NRQM [10]
 $(1.54\pm 0.40)$$\times~{}10^{-10}$ EFT+LCDA [103]
$\rm B^{*\pm}$  $(1.3\pm 0.2)$$\times~{}10^{-5}$ EFT+NRQM [10]
 $(1.41\pm 0.36)$$\times~{}10^{-10}$ EFT+LCDA [103]
$\rm B_{c}^{\pm}$  $(1.6\pm 0.2)$$\times~{}10^{-8}$ EFT+NRQM [10]
 $(8.21\pm 0.83)$$\times~{}10^{-8}$ EFT+LCDA [103]
@@ III.3 Radiative Higgs leptonium decays
Figure 9 shows the diagrams of the decay of the Higgs boson into a photon or a Z boson plus a leptonium state $(\ell^{+}\ell^{-})$, where $(\ell^{+}\ell^{-})=(\mathrm{e}^{+}\mathrm{e}^{-}),\,(\mu^{+}\mu^{-}),\,(\tau^{+}\tau^{-})$ represent positronium, dimuonium, and ditauonium bound states, respectively. Depending on the accompanying gauge boson (Z or $\gamma$), leptonium can be produced in spin triplet (ortho) and/or spin singlet (para) states. Such decays are similar to the exclusive radiative decays into mesons shown in Fig. 4, but changing the quark lines for lepton lines. Of course, in the SM there is no unlike-flavour leptonic decay possible, at variance with the $\rm H\to W^{\pm}+M$ case. However, in the presence of BSM physics leading to LFV Higgs decays, $(\ell\ell^{\prime})$ bound states could be formed directly in exclusive radiative decays. We provide here an estimate of the decay rates for the SM processes by considering the similar $\rm H\to\gamma+M,\,Z+M$ processes, and changing the quarkonium form-factors by leptonium ones. If such decays were to be produced with relatively large rates, they would provide interesting cross checks for any LFUV effects potentially observed with the “open” leptons [116].
The overall probability for forming an onium bound state is determined by one single parameter: its radial wavefunction at the origin of the space coordinate, which for leptonium bound states (of principal quantum number $n$) reads [117],
@@
Since the QED coupling is much smaller than the QCD one, $\alpha(m_{\ell\ell})\ll C_{F}\alpha_{\text{s}}(m_{\mathrm{q}\overline{\mathrm{q}}})$, and since the charged lepton masses are smaller than the quark masses of the same generation, $m_{\ell\ell}\ll m_{\mathrm{q}\overline{\mathrm{q}}}$, the ratio $[\alpha(0)m_{\ell\ell}/(\alpha_{\mathrm{s}}(m_{\mathrm{q}\overline{\mathrm{q}}})m_{\mathrm{q}\overline{\mathrm{q}}})]^{3}$ is very small, and one can anticipate that those decays will be orders-of-magnitude more suppressed than the $\rm H\to\gamma+M(\mathrm{q}\overline{\mathrm{q}})$ ones discussed in Section III.2.
@@
We start by calculating the branching fraction for the $\rm H\to\gamma+(\ell^{+}\ell^{-})$ process in which, due to charge-conjugation conservation, the leptonium can only be in the ortho $(\ell^{+}\ell^{-})_{1}$ state. The partial width of this decay can be derived from the similar expressions for quarkonium vector mesons [11], namely
@@
where for the second equality of the indirect component, we have adopted the Van Royen–Weisskopf formula [118, 119] to relate the leptonium decay constant, $f_{\mathrm{M}}$, to its wavefunction at the origin:
@@
Here, the particle M can be either a pseudoscalar/vector meson or a para/ortho-leptonium state, and the number of colours $N_{\mathrm{c}}=1$ is used for the latter. We choose $\phi_{\mathrm{M}}(0)$ to be real, and then the decay constant $f_{\mathrm{M}}$ comes with a phase that decides the interference between indirect and direct amplitudes. In this work, we choose the (positive) phase that yields a destructive interference similar to the quarkonium case [11]. Since $m_{\ell\ell}\ll m_{\mathrm{H}}$, plugging Eq. (24) into (27), one can see that the indirect amplitudes are independent of the leptonium masses. Using Eqs. (25)–(27) with the numerical values of Table 2, we determine the radiative ortholeptonium $(n=1)$ branching fractions listed in the first three rows of Table 8. The radiative ortholeptonium decays of the Higgs boson have all numerically similar $\mathcal{O}(10^{-12})$ rates for the three leptons, they have not been searched-for to date at the LHC, and only a future machine such as the FCC-hh can try to set upper limits on them at about 10 times their SM values. The experimental search has a very clean signature characterized by a secondary vertex from the boosted $(\ell^{+}\ell^{-})_{1}$ decay, which leads to significantly displaced triphotons (for positronium and dimuonium), $\mathrm{e}^{+}\mathrm{e}^{-}$ pairs (for dimuonium and ditauonium), or $\mu^{+}\mu^{-}$ pairs (for ditauonium).
We determine next the rates of the $\rm H\to Z+(\ell^{+}\ell^{-})$ decays. At variance with the photon case, these decays have a massive final-state gauge boson that can be in a longitudinal polarization state. As a consequence, both scalar (para-) and vector (ortho-) leptonium states can be produced, whereas in the $\rm H\to\gamma+(\ell^{+}\ell^{-})_{1}$ case, the leptonium could only be a (transversely polarized) ortho state. We consider first the Higgs decays into Z-boson-plus-ortholeptonium by properly adapting the theoretical expressions for the similar $\rm H\to Z+VM$ decays, where VM are quarkonia vector mesons such as $\mathrm{J}/\psi,\Upsilon$ [108]. The corresponding $\rm Z+(\ell^{+}\ell^{-})_{1}$ decay width can then be written as
@@
where $\Gamma_{1},\,\Gamma_{2},$ and $\Gamma_{3}$ are the contributions from $\rm H\to Z+Z^{*}\to Z+(\ell^{+}\ell^{-})_{1}$, $\rm H\to Z+\gamma^{*}\to Z+(\ell^{+}\ell^{-})_{1}$, and their interference, respectively. Using $r_{\mathrm{Z}}=m_{\mathrm{Z}}^{2}/m_{\mathrm{H}}^{2}$, $r_{\ell\ell}=m_{\ell\ell}^{2}/m_{\mathrm{H}}^{2}$, the Källén function $\lambda(a,b,c)=a^{2}+b^{2}+c^{2}-2(ab+ac+bc)$, and the $\rm H\gamma Z$ effective coupling $C_{Z\gamma}\approx 5.54$ [120], these individual partial widths can be expressed as follows
@@
Lastly, we compute the rates of the $\rm H\to Z+(\ell^{+}\ell^{-})_{0}$ decays, whose width can be written using similar expressions for the $\rm H\to Z+M$ decays (where M is a pseudoscalar meson) [103], as follows
@@
with $a_{\mathrm{\ell}}=T_{3}^{\mathrm{\ell}}/2=1/4$, which is suppressed by a factor $m_{\ell\ell}^{2}/m_{\mathrm{H}}^{2}$ or $m_{\mathrm{\ell}}^{2}/m_{\mathrm{H}}^{2}$, that makes it completely negligible (the contribution to the final $(\tau\tau)_{0}$ amplitude is of about 0.09%). The contribution from the indirect amplitude to the width (31) reads

### v4

@@ III.2.3 Higgs decays into a gauge boson plus a flavoured meson
On the one hand, calculations for the inclusive flavour-changing $\rm H\to qq^{\prime}$ rates exist [Benitez-Guzman:2015ana, Aranda:2020tqw, Kamenik:2023hvi], but without the photon (or Z boson) emission and the exclusive final-state meson formation. These predicted FCNC Higgs branching fractions, $\mathcal{B}(\rm H\to qq^{\prime})\equiv\mathcal{B}(\rm H\to\mathrm{q}\overline{\mathrm{q}}^{\prime}+q^{\prime}\overline{q})$, amount to  [^5] The $\rm H\to ds$ branching fraction has been updated as per Ref. [Tammaro2025].
@@
$\mathcal{B}(\rm H\to sb)$ $=$ $(8.9\pm 1.5)\cdot 10^{-8},\;\,\quad\mathcal{B}(\rm H\to ds)=(1.9\pm 0.3)\cdot 10^{-15}.$ (17)
On the other hand, the work of [Kagan:2014ila] determined the $\rm H\to\gamma+VM$ branching fractions, and the work [Alte:2016yuw] provided the $\rm H\to Z+VM$ decay widths assuming (arbitrary) $\mathcal{O}(1)$ flavour-changing Yukawa couplings. From the results of [Kagan:2014ila] (resp., [Alte:2016yuw]), the branching ratios of Higgs decaying into a photon (resp., a Z boson) plus a flavoured neutral meson can be determined through the following EFT + LCDA-based expressions
@@
where $p,\ \lambda$ are the momentum and polarization of the vector meson, respectively, and $\mu$ is the factorization scale. The corresponding numerical values used here are [Pullin:2021ebn, RBC-UKQCD:2008mhs]: $f^{\perp}_{\rm B^{*0}}(3\text{ GeV})=0.2\text{ GeV}$, $f^{\perp}_{\rm B^{*0}_{\rm s}}(3\text{ GeV})=0.236\text{ GeV}$, $f^{\perp}_{\rm K^{*0}}(2\text{ GeV})=0.717f_{\rm K^{*0}}$, $f^{\perp}_{\rm D^{*0}}(2\text{ GeV})=0.202\text{ GeV}$. These transverse decay constants can be evolved to the higher scale $\mu_{\rm HZ}$ using the renormalization group equation [Konig:2015qat], thereby giving: $f^{\perp}_{\rm VM}(\mu_{\rm HZ})=0.86\ f^{\perp}_{\rm VM}(3\text{ GeV})=0.83\ f^{\perp}_{\rm VM}(2\text{ GeV})$, using $\alpha_{\mathrm{s}}(\mu_{\rm HZ})=0.126$, $\alpha_{\mathrm{s}}(3\text{ GeV})=0.25$, $\alpha_{\mathrm{s}}(2\text{ GeV})=0.29$, and $N_{\mathrm{f}}=5$ number of flavours.
@@
$\mathcal{L}_{H\to\overline{q}q^{\prime}}=\sum_{q}\kappa_{\rm qq^{\prime}}h\overline{q}_{L}q^{\prime}_{R}+\text{h.c.}\ ,$ (21)
@@
$\begin{cases}|\kappa_{\rm bs}|^{2}+|\kappa_{\rm sb}|^{2}=2.4\times 10^{-11},\\ |\kappa_{\rm bd}|^{2}+|\kappa_{\rm db}|^{2}=1.0\times 10^{-12},\\ |\kappa_{\rm cu}|^{2}+|\kappa_{\rm uc}|^{2}=7.4\times 10^{-24},\\ |\kappa_{\rm ds}|^{2}+|\kappa_{\rm sd}|^{2}=5.2\times 10^{-19}.\end{cases}$ (23)
Combining Eq. (23) with (18) or (19), we arrive at the theoretical branching fractions listed in Table 6. Such exclusive FCNC radiative decays are extremely suppressed in the SM, in the range of $\mathcal{O}(10^{-14}$$10^{-30})$. Therefore, radiative flavoured meson decays of the Higgs boson appear utterly rare to be visible at any current or future collider, and therefore a very clean probe of any BSM physics that may enhance Higgs FCNC decays. Experimental limits have been recently set for the $\rm H\to\gamma+K^{*0},\gamma+D^{*0}$ channels at $\mathcal{O}(10^{-4}\mbox{--}10^{-3})$ [ATLAS:2023alf, ATLAS:2024dpw], to be compared with our $\mathcal{O}(10^{-23}\mbox{--}10^{-27})$ SM predictions.
Table 6: Compilation of exclusive Higgs decays to a gauge boson plus a flavoured meson. For each decay, we provide the branching fraction predicted using Eqs. (18), (19), as well as the current experimental upper limit and that estimated for HL-LHC. The last two columns indicate whether the decay can be produced at FCC-ee/FCC-$h$.
@@
$\rm H\to\,X$ ++ $\rm M$ Branching fraction Framework 2024 HL-LHC FCC-$e$ FCC-$h$
[…]
@@
$\rm H\,\to\,Z$ + $\rm K^{*0}$ $2.2$ $\times~10^{-25}$ EFT+LCDA (this work)
@@ III.2.4 Higgs decays into a W boson plus a meson
The charged $\rm H\to W^{\pm}M^{\mp}$ decays (corresponding to the diagrams of Fig. 4 with W bosons in the final state) differ qualitatively from the neutral radiative decays discussed above, because the W attaches itself to a charged current, and one can probe flavour-violating couplings of the Higgs boson. The theoretical complication is that the W has both transverse and longitudinal polarizations, yielding lengthier analytical expressions [Isidori:2013cla, Alte:2016yuw]. Table 7 lists the theoretical predictions for Higgs decays into a W± boson plus a charged meson, and Fig. 8 presents them in graphical form. The rates for these processes range roughly between $10^{-5}$ and $10^{-10}$ with small differences between models for the same final state. The EFT + NRQM theoretical predictions of Ref. [LHCHiggsCrossSectionWorkingGroup:2016ypw], which update the $\mathcal{B}$ numerical values computed in [Isidori:2013cla] (but have likely a typo in the $\mathcal{B}(\rm H\to B^{*\pm}W^{\mp})=10^{-5}$ rate quoted, which should be $10^{-10}$), agree with the alternative EFT + LCDA rates of Ref. [Alte:2016yuw]. There has been no experimental search to date at the LHC for any of these 11 decays. The most promising channel, with a $\mathcal{O}(10^{-5})$ branching fraction, is $\rm H\to W^{\pm}+\rho^{\mp}$ with the rho meson decaying into two pions with almost 100% probability, but seemingly experimental evidence for 7 (all) of them will only be possible at FCC-ee (FCC-hh).
Table 7: Compilation of exclusive Higgs decays to a W boson plus a meson. For each decay, we provide the predicted branching fraction(s) and the theoretical approach used to compute it. The last two columns indicate whether the decay can be produced at FCC-ee/FCC-$h$.
@@
$\rm H\to\,W$ ++ $\rm M$ Branching fraction Framework 2024 HL-LHC FCC-$e$ FCC-$h$
$\rm H\,\to\,W^{\mp}$ + $\rm\pi^{\pm}$ $(4.2\pm 0.2)$ $\times~10^{-6}$ EFT+NRQM [LHCHiggsCrossSectionWorkingGroup:2016ypw]
$(4.3\pm 0.2)$ $\times~10^{-6}$ EFT+LCDA [Alte:2016yuw]
$\rm\rho^{\pm}$ $(1.5\pm 0.1)$ $\times~10^{-5}$ EFT+NRQM [LHCHiggsCrossSectionWorkingGroup:2016ypw]
$(1.09\pm 0.05)$ $\times~10^{-5}$ EFT+LCDA [Alte:2016yuw]
$\rm K^{\pm}$ $(3.3\pm 0.1)$ $\times~10^{-7}$ EFT+NRQM [LHCHiggsCrossSectionWorkingGroup:2016ypw]
$(3.3\pm 0.1)$ $\times~10^{-7}$ EFT+LCDA [Alte:2016yuw]
$\rm K^{*\pm}$ $(4.3\pm 0.2)$ $\times~10^{-7}$ EFT+NRQM [LHCHiggsCrossSectionWorkingGroup:2016ypw]
$(5.6\pm 0.4)$ $\times~10^{-7}$ EFT+LCDA [Alte:2016yuw]
$\rm D^{\pm}$ $(5.8\pm 0.6)$ $\times~10^{-7}$ EFT+NRQM [LHCHiggsCrossSectionWorkingGroup:2016ypw]
$(5.6\pm 0.5)$ $\times~10^{-7}$ EFT+LCDA [Alte:2016yuw]
$\rm D^{*\pm}$ $(1.3\pm 0.1)$ $\times~10^{-6}$ EFT+NRQM [LHCHiggsCrossSectionWorkingGroup:2016ypw]
$(1.04\pm 0.14)$ $\times~10^{-6}$ EFT+LCDA [Alte:2016yuw]
$\rm D_{s}^{\pm}$ $(1.6\pm 0.1)$ $\times~10^{-5}$ EFT+NRQM [LHCHiggsCrossSectionWorkingGroup:2016ypw]
$(1.71\pm 0.11)$ $\times~10^{-5}$ EFT+LCDA [Alte:2016yuw]
$\rm D_{s}^{*\pm}$ $(3.5\pm 0.2)$ $\times~10^{-5}$ EFT+NRQM [LHCHiggsCrossSectionWorkingGroup:2016ypw]
$(2.51\pm 0.19)$ $\times~10^{-5}$ EFT+LCDA [Alte:2016yuw]
$\rm B^{\pm}$ $(1.6\pm 0.4)$ $\times~10^{-10}$ EFT+NRQM [LHCHiggsCrossSectionWorkingGroup:2016ypw]
$(1.54\pm 0.40)$ $\times~10^{-10}$ EFT+LCDA [Alte:2016yuw]
$\rm B^{*\pm}$ $(1.3\pm 0.2)$ $\times~10^{-5}$ EFT+NRQM [LHCHiggsCrossSectionWorkingGroup:2016ypw]
$(1.41\pm 0.36)$ $\times~10^{-10}$ EFT+LCDA [Alte:2016yuw]
$\rm B_{c}^{\pm}$ $(1.6\pm 0.2)$ $\times~10^{-8}$ EFT+NRQM [LHCHiggsCrossSectionWorkingGroup:2016ypw]
$(8.21\pm 0.83)$ $\times~10^{-8}$ EFT+LCDA [Alte:2016yuw]
@@ III.3 Radiative Higgs leptonium decays
Figure 9 shows the diagrams of the decay of the Higgs boson into a photon or a Z boson plus a leptonium state $(\ell^{+}\ell^{-})$, where $(\ell^{+}\ell^{-})=(\mathrm{e}^{+}\mathrm{e}^{-}),\,(\mu^{+}\mu^{-}),\,(\tau^{+}\tau^{-})$ represent positronium, dimuonium, and ditauonium bound states, respectively. Depending on the accompanying gauge boson (Z or $\gamma$), leptonium can be produced in spin triplet (ortho) and/or spin singlet (para) states. Such decays are similar to the exclusive radiative decays into mesons shown in Fig. 4, but changing the quark lines for lepton lines. Of course, in the SM there is no unlike-flavour leptonic decay possible, at variance with the $\rm H\to W^{\pm}+M$ case. However, in the presence of BSM physics leading to LFV Higgs decays, $(\ell\ell^{\prime})$ bound states could be formed directly in exclusive radiative decays. We provide here an estimate of the decay rates for the SM processes by considering the similar $\rm H\to\gamma+M,\,Z+M$ processes, and changing the quarkonium form-factors by leptonium ones. Apart from probing lepton Yukawa couplings, if such decays were to be produced with relatively large rates, they would provide interesting cross checks for any LFUV effects potentially observed with the “open” leptons [Fael:2018ktm].
The overall probability for forming an onium bound state is determined by one single parameter: its radial wavefunction at the origin of the space coordinate, which for leptonium bound states (of principal quantum number $n$) reads [dEnterria:2022alo],
@@
Since the QED coupling is much smaller than the QCD one, $\alpha(m_{\ell\ell})\ll C_{F}\alpha_{\text{s}}(m_{\mathrm{q}\overline{\mathrm{q}}})$, and since the charged lepton masses are smaller than the quark masses of the same generation, $m_{\ell\ell}\ll m_{\mathrm{q}\overline{\mathrm{q}}}$, the ratio $\alpha(0)m_{\ell\ell}/(\alpha_{\mathrm{s}}(m_{\mathrm{q}\overline{\mathrm{q}}})m_{\mathrm{q}\overline{\mathrm{q}}})]^{3}$ is very small, and one can anticipate that those decays will be orders-of-magnitude more suppressed than the $\rm H\to\gamma+M(\mathrm{q}\overline{\mathrm{q}})$ ones discussed in Section III.2.
@@
We start by calculating the branching fraction for the $\rm H\to\gamma+(\ell^{+}\ell^{-})$ process in which, due to charge-conjugation conservation, the leptonium can only be in the ortho $(\ell^{+}\ell^{-})_{1}$ state. The partial width of this decay can be derived from the similar expressions for quarkonium vector mesons [Bodwin:2013gca], namely
@@
where for the second equality of the indirect component, we have adopted the Van Royen–Weisskopf formula [Azhothkaran:2020ipl, VanRoyen:1967nq] to relate the leptonium decay constant, $f_{\mathrm{M}}$, to its wavefunction at the origin:
@@
Here, the particle M can be either a pseudoscalar/vector meson or a para/ortho-leptonium state, and the number of colours $N_{\mathrm{c}}=1$ is used for the latter. We choose $\phi_{\mathrm{M}}(0)$ to be real, and then the decay constant $f_{\mathrm{M}}$ comes with a phase that decides the interference between indirect and direct amplitudes. In this work, we choose the (positive) phase that yields a destructive interference similar to the quarkonium case [Bodwin:2013gca]. Since $m_{\ell\ell}\ll m_{\mathrm{H}}$, plugging Eq. (24) into (27), one can see that the indirect amplitudes are independent of the leptonium masses. Using Eqs. (25)–(27) with the numerical values of Table 2, we determine the radiative ortholeptonium $(n=1)$ branching fractions listed in the first three rows of Table 8. The radiative ortholeptonium decays of the Higgs boson have all numerically similar $\mathcal{O}(10^{-12})$ rates for the three leptons, they have not been searched-for to date at the LHC, and only a future machine such as the FCC-hh can try to set upper limits on them at about 10 times their SM values. The experimental search has a very clean signature characterized by a secondary vertex from the boosted $(\ell^{+}\ell^{-})_{1}$ decay, which leads to significantly displaced triphotons (for positronium and dimuonium), $\mathrm{e}^{+}\mathrm{e}^{-}$ pairs (for dimuonium and ditauonium), or $\mu^{+}\mu^{-}$ pairs (for ditauonium).
We determine next the rates of the $\rm H\to Z+(\ell^{+}\ell^{-})$ decays. At variance with the photon case, these decays have a massive final-state gauge boson that can be in a longitudinal polarization state. As a consequence, both scalar (para-) and vector (ortho-) leptonium states can be produced, whereas in the $\rm H\to\gamma+(\ell^{+}\ell^{-})_{1}$ case, the leptonium could only be a (transversely polarized) ortho state. We consider first the Higgs decays into Z-boson-plus-ortholeptonium by properly adapting the theoretical expressions for the similar $\rm H\to Z+VM$ decays, where VM are quarkonia vector mesons such as $\mathrm{J}/\psi,\Upsilon$ [Gao:2014xlv]. The corresponding $\rm Z+(\ell^{+}\ell^{-})_{1}$ decay width can then be written as
@@
where $\Gamma_{1},\,\Gamma_{2},$ and $\Gamma_{3}$ are the contributions from $\rm H\to Z+Z^{*}\to Z+(\ell^{+}\ell^{-})_{1}$, $\rm H\to Z+\gamma^{*}\to Z+(\ell^{+}\ell^{-})_{1}$, and their interference, respectively. Using $r_{\mathrm{Z}}=m_{\mathrm{Z}}^{2}/m_{\mathrm{H}}^{2}$, $r_{\ell\ell}=m_{\ell\ell}^{2}/m_{\mathrm{H}}^{2}$, the Källén function $\lambda(a,b,c)=a^{2}+b^{2}+c^{2}-2(ab+ac+bc)$, and the $\rm H\gamma Z$ effective coupling $C_{Z\gamma}\approx 5.54$ [Korchin:2013ifa], these individual partial widths can be expressed as follows
@@
Lastly, we compute the rates of the $\rm H\to Z+(\ell^{+}\ell^{-})_{0}$ decays, whose width can be written using similar expressions for the $\rm H\to Z+M$ decays (where M is a pseudoscalar meson) [Alte:2016yuw], as follows
@@
with $a_{\mathrm{\ell}}=T_{3}^{\mathrm{\ell}}/2=1/4$, which is suppressed by a factor $m_{\ell\ell}^{2}/m_{\mathrm{H}}^{2}$, that makes it completely negligible (the contribution to the final $(\tau\tau)_{0}$ amplitude is of about 0.09%). The contribution from the indirect amplitude to the width (31) reads